FOURIER TRANSFORM EMISSION SPECTROSCOPY OF THE g^{4}Δ − a^{4}Δ SYSTEM OF FeCl

\underline{R. S. RAM}, Department of Chemistry, University of Arizona, Tucson, AZ 85721; and P. F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ont., Canada N2L 3G1.

The emission spectrum of FeCl has been investigated at high resolution in the near infrared. The molecule was excited in a microwave discharge lamp using a mixture of FeCl_{3} vapor and 2.5 Torr of He, and the spectra were recorded in the 3000-12500 cm^{−1} region using a Fourier transform spectrometer. New bands with R heads near 7725, 8149, 8577, 8949, 9319, 9686 and 10050 cm^{−1} have been assigned as the 0-2, 0-1, 0-0, 1-0, 2-0, 3-0 and 4-0 bands of a new g^{4}Δ−a^{4}Δ system of FeCl. The strong bands of this system consist of four subbands assigned as ^{4}Δ_{1/2}-^{4}Δ_{1/2}, ^{4}Δ_{3/2}-^{4}Δ_{3/2}, ^{4}Δ_{5/2}-^{4}Δ_{5/2} and ^{4}Δ_{7/2}-^{4}Δ_{7/2}. This transition is analogous to the 1 μm transition of FeH and FeF. A rotational analysis of a number of bands in each subband has been obtained and spectroscopic constants extracted. The lower a^{4}Δ state is affected by perturbations from a nearby electronic state while the excited g^{4}Δ state is free from perturbation. This work represents the first observation of a transition involving the quartet states of FeCl.